OpenStudy (anonymous):

cos^2x-sin^2x=1-2sin^2x need help showing step by step to prove identity

OpenStudy (campbell_st):

well you know cos^2 + sin^2 = 1 make cos^2 the substitute and then substitute into the equation...

OpenStudy (jdoe0001):

\(\bf cos^2(\theta)+sin^2(\theta)=1\implies cos^2(\theta)={\color{blue}{ 1-sin^2(\theta)}}\\ \quad \\ \quad \\ cos^2(x)-sin^2(x)\implies {\color{blue}{ 1-sin^2(x)}}-sin^2(x)\)